Q.1 How many ways can 4 teddy bears , 3 toy trucks , 5 dolls and 1 toy cars be distributed among 13 children if each child is to receive 1 item?
Q.2 In a group of 14 students there are 8 girls and 6 boys.
The committee of 4 students must have 1 girl and 1 boy and 2 other students, how many different committees can be chosen from this group?

Answer by stanbon(75887)

Q.1 How many ways can 4 teddy bears , 3 toy trucks , 5 dolls and 1 toy cars be distributed among 13 children if each child is to receive 1 item?
Ans:: 13! = 62270208.. ways
=================
Q.2 In a group of 14 students there are 8 girls and 6 boys.
The committee of 4 students must have 1 girl and 1 boy and 2 other students, how many different committees can be chosen from this group?
Ans: 8*6*12C2 = 48*66 = 3168 different committees
--------------
Cheers,
Stan H.
------------